RC Meeting, Thursday, September 10, 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meeting at 16:30, (remote)

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • Taking the beam we can get!

Current Run Plan

  • Finished the the LH2 and dummy target runs for the first part of the first setting
  • Tweaks to the analysis have cleaned things up a bit, we can cut on the coin time and get a nice MM spectrum, various detector calibrations need fixing
    • The hodoscope particularly needs work
    • See here for some nice plots!
  • Running LD2, unfortunately, need to redo an hour or so of data (see Hall status)
  • If things go well, we will need to rotate either late Owl or early Day tomorrow
    • An extremely optimistic scenario is late Swing, early Owl today

Planned Accesses to the Hall

  • Joe B and Larry both requested an access when possible
    • See Hall Status, will drop to controlled during Hall D setup
    • Scheduled to start Monday 06:00 - Let me know if you can do your work with an escorted access or not

Accelerator Status

  • Very smooth through most of swing yesterday, issue with the north linac took us down from ~ 22:00 to 00:30
  • Beam a little trippy overnight, OK so far today

Hall Status

  • CODA a little unstable - rates too high
    • For Physics settings, reduce the beam current, do not change the coin prescale (PS5) from 0
  • COIN rates on the LD2 are extremely high, this is what prompted the Owl shift workers to try adjusting the prescales (they were seeing enormous deadtimes and rates - ~ 7kHz un-prescaled coin rate at 70 uA)
  • This will only get worse at smaller angles
  • Beam trips causing an alarm when running with LD2 - related to the fan?

Notices in Effect for Shift Crew

Plans going forward

  • BCM calibrations, do over the weekend?
  • Arc energy measurement, over the weekend if possible
  • Bullseye scan